OpenStudy (anonymous):

Simplify: (15mn^2p)/25m^2n^2

OpenStudy (anonymous):

3p/5m

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(3*5*n^2 *m*p)/(5*5*m^2 *n^2)= cancel out: 5 & 5 (dominator & numerator): 3/5 * n^(2-2) * p * m(1-2)=3/5 * p * m^(-1) or: (3p)/(5m)